Titanium reacts with iodine to form titanium(III) iodide, emitting heat. 2Ti(s)+3I2(g)→2TiI3(s)ΔHorxn = -839 kJ

Determine the masses of titanium and iodine that react if 1.40×103 kJ of heat is emitted by the reaction.

Express your answer to three significant figures and include the appropriate units
Answer #1

The reaction:

2Ti(s)+3I2(g)→2TiI3(s)ΔHorxn = -839 kJ

Heat requried =  1.40×103 kJ

ratio -- >= (1.4*10^3)/(839) = 1.66865

then

mol of Ti required = 2*1.66865 = 3.3373

mol of I2 = 3*1.66865 = 5.00595

now...

mass of Ti = mol*MW = 3.3373*47.867 = 159.7465 g of Ti

mass of I2 = mol*MW = 5.00595*253.809 = 1270.55 g of I2